OpenStudy (anonymous):

WILL GIVE MEDAL Given a point and a slope, write the equation of the line in point- slope form then change into slope-intercept form m=5/6,(7,3) PLEASE SHOW WORK

OpenStudy (anonymous):

first step is to write \[y-3=\frac{5}{6}(x-7)\]

OpenStudy (anonymous):

second step is to distribute the \(\frac{5}{6}\) on the right to get \[y-3=\frac{5}{6}x-\frac{35}{6}\]

OpenStudy (anonymous):

third step is to add \(3\) to both sides to get \(y\) by itself

OpenStudy (anonymous):

I'm bad at adding fractions...

OpenStudy (anonymous):

\[3=\frac{18}{6}\] \[\frac{18}{6}-\frac{35}{6}=-\frac{17}{6}\]

OpenStudy (anonymous):

everyone is bad at adding fractions